Q2

if two events, A and B, are such that P(A) = .5, P(B) = .3, and P(A intersection B) = .1, findthe following:

a).P(A|B)

b). P(B|A)

c).P(A|AUB)

d).P(A|A intersection B)

e). P(A intersection B|A U B)


1
Expert's answer
2021-06-28T08:53:03-0400

a).

"P(A|B)=\frac{P(A\bigcap B)}{P(B)}\\
P(A\bigcap B)=0.1\\
P(B)=0.3\\
P(A|B)=\frac{0.1}{0.3}=\frac{1}{3}\\"

b).

"P(B|A)=\frac{P(B\bigcap A)}{P(A)}\\
P(B\bigcap A)=P(A\bigcap B)=0.1\\
P(A)=0.5\\
P(B|A)=\frac{0.1}{0.5}=\frac{1}{5}\\"

c).

"P(A|A\bigcup B)=\frac{P(A\bigcap (A\bigcup B))}{P(A\bigcup B)}\\
P(A|A\bigcup B)=\frac{P(A)}{P(A\bigcup B)}\\
P(A)=0.5\\
P(A\bigcup B)=P(A)+P(B)-P(A\bigcap B)\\
P(A\bigcup B)=0.5+0.3-0.1=0.7\\
P(A|A\bigcup B)=\frac{0.5}{0.7}=\frac{5}{7}\\"

d).

"P(A|A\bigcap B)=\frac{P(A\bigcap (A\bigcap B))}{P(A\bigcap B)}\\
P(A|A\bigcap B)=\frac{P(A\bigcap B)}{P(A\bigcap B)}=1\\"

e).

"P(A\bigcap B|A \bigcup B)=\frac{P(A\bigcap B\bigcap A \bigcup B)}{P(A \bigcup B)}\\
P(A\bigcap B|A \bigcup B)=\frac{P(A\bigcap B)}{P(A \bigcup B)}\\
P(A\bigcap B|A \bigcup B)=\frac{0.1}{0.7}=\frac{1}{7}\\"